OpenStudy (anonymous):

Solve for x: 2 over quantity x minus 2 plus 7 over quantity x squared minus 4 equals 5 over x

OpenStudy (anonymous):

\[\frac{2}{x-2}+\frac{7}{x^2-4}=\frac{5}{x}~~?\]

OpenStudy (anonymous):